OpenStudy (anonymous):

simplify: 3 square root of 45 minus 2 square root of 125

OpenStudy (anonymous):

\[3\sqrt{45}-2\sqrt{125}\]

OpenStudy (anonymous):

and simplify this \[(-27)^{-1/3}\]

OpenStudy (anonymous):

\[3\sqrt{9*5}-2\sqrt{25*5}\] \[3*3\sqrt{5}-2*5\sqrt{5}\] \[9\sqrt{5}-10\sqrt{5}\]= \[-1\sqrt{?}\]

OpenStudy (anonymous):

thats is suppose to be \[-1\sqrt{5}\]

OpenStudy (anonymous):

and have you tried simplifying that other one?

OpenStudy (anonymous):

not yet.

OpenStudy (anonymous):

can you simplify it?

OpenStudy (anonymous):

sure i can

OpenStudy (anonymous):

Please and thank you, sir :)

OpenStudy (anonymous):

first of all realize that we can write this problem as : \[1/\sqrt[3]{-27}\]

OpenStudy (anonymous):

and answer is -1/3

OpenStudy (anonymous):

thank you :)